Circular No. 5971

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1994J IN ANONYMOUS GALAXY C. Pollas, Observatoire de la Côte d’Azur, reports his discovery of a

supernova of mag 19 on Technical Pan films taken with the OCA 0.9-m Schmidt telescope on Mar. 5.97 and 8.91 UT by D. Albanese and himself. The supernova is located at α = 9^{h}58^{m}15^{s}.82, δ = +54^{o}49′24′′.0 (equinox 1950.0), which is 2′′.7 east and 3′′.4 south of the center of a spiral galaxy of mag 17. Nothing appears on a similar film obtained on 1994 Jan. 18.1 to limiting mag 22. A nearby star of mag ∼ 18 is at (end figures) 10^{s}.52, 48′45′′.7. A. V. Filippenko, A. J. Barth, and T. Matheson, University of California at Berkeley, report that fully calibrated CCD spectra (range 425-702 nm) obtained on Apr. 4 UT with the Shane 3-m reflector at Lick Observatory confirm that this object is a supernova, but the spectral type is uncertain due to the low S/N ratio. The redshift of the parent galaxy is 0.056, based on measurements of several emission lines.

SUPERNOVA 1994I IN NGC 5194 M. Turatto and C. Zanin, Padova, communicate: “We observed SN

1994I on Apr. 9.0 UT with the Cima Ekar 1.8-m telescope (+ Boller & Chivens spectrograph + CCD) under poor weather conditions. The preliminary reduction of the spectrogram (range 350-800 nm, resolution 2 nm) shows a number of broad emissions at 367, 393, 460, 488, 535, 592, and 650 nm and absorptions at 374, 437, 478, 498, 570, and 619 nm. The spectrum shows interesting resemblance to some peculiar type-Ia supernovae shortly before maximum. In particular, the spectrum of SN 1994I blueward of 480 nm closely matches those of SNe 1986G and 1991bg. The absorption at 619 nm may be due to Si II. However, between 500 and 600 nm, the observed spectrum is dominated by a very broad emission feature and differs from those of SNe 1986G and 1991bg.”

OJ 287 H. Meusinger, Thüringer Landessternwarte, Tautenburg, reports the

following photographic B magnitudes of OJ 287 (cf. IAUC 5909) from Tautenburg Schmidt plates (emulsion ORWO ZU, filter GG 13): Feb. 14.9 UT, 16.1; Mar. 10.9, 16.0; Apr. 5.8, 15.2.

1994 April 9 Daniel W. E. Green